Example of the Abstract for International Workshop on Vortex Physics in High-T_{c} Superconductrs

Kazuto Hirata

National Research Institute for Metals, Tsukuba 305-0047, Japan

Abstract

This is an example of the abstract. Authors can describe their abstarcts within about 21 lines in the length of lines in this text. This corresponds to one page of A4 size except three lines of the title, the authors’ name and the address. If there are many authors, the text must be shortened. It will be a great help for us if you write the abstarct in REVTEX style. Followings are typical examples of Greek, subscript, superscript: Bi_{2}Sr_{2}CaCu_{2}O_{8+δ}, B_{Φ}, ω_{pl} = \sqrt{8π^{2}csj_{c}^{c}/ϵ_{0}Φ_{0}}, λ_{ab}(T) = λ_{ab}(0)/\sqrt{1 − (T/T_{c})^{2}} . Formula can also be expressed like,

ω_{pl}^{2}(B, T) = ω_{pl}^{2}(0, T) < cos φ_{n,n+1} > .  (1)

If authors want to put the references in the text, please put them in the text like; It has been confirmed [Zeldovet al., Phys.Rev.Lett.72,1914 (1994)].